OpenStudy (anonymous):

can this be differentiated? And if so how. y = sqrt(x) - (1/2)^x

OpenStudy (jamesj):

For any positive number \( a \), suppose \( y = a^x \). Then \[ \ln y = x \ln a \] and hence \[ \frac{1}{y}y' = \ln a \] \[ y' = y \ln a = (\ln a)a^x \] Therefore for the function \( f(x) = (1/2)^x \), the derivative is \[ \frac{df}{dx} = \ln(1/2).(1/2)^x = -(\ln 2) 2^{-x} \]

OpenStudy (jamesj):

Make sense?